i have to find the midpoint of the line between the two given points....
1. (10,4)and(2,6) 2. (-2,-1)and (-8,6)

Found 2 solutions by stanbon, Nate:
Answer by stanbon(75887)

i have to find the midpoint of the line between the two given points....
1. (10,4)and(2,6)
The midpoint coordinates are the average of the given coordinates.
x-coordinate = (10+2)/2=6
y-coordinate = (4+6)/2 =5
midpoint is (6,5)
---------------------------
2. (-2,-1)and (-8,6)
x-coordinate = (-2-8)/2 = -5
y-coordinate = (-1+6)/2 = 5/2
Midpoint is (-5,5/2)
------------------
Cheers,
Stan H.

Answer by Nate(3500)

1. (10,4)and(2,6)
You sum the two and divide by 2.
(10 + 2)/2 = 12/2 = 6
(4 + 6)/2 = 10/2 = 5
(6,5)
It is finding the average. I'll let you solve the next.